# On the convergence of loss and uncertainty-based active learning algorithms

### v2

@@ 5 Conclusion
In this paper, we have provided convergence rate guarantees for both loss and uncertainty-based active learning algorithms under various assumptions. Furthermore, we introduced a novel algorithm that combines sampling with an adaptive size, conforming to stochastic Polyak’s step size in expectation, and demonstrated its convergence rate guarantee, contingent on a condition related to the sampling probability function. For future research, it would be interesting to apply the methodologies presented in this paper to obtain additional convergence rate results. Additionally, exploring a theoretical study on the impact of noise in the loss estimator, used for evaluating the sampling probability function, on the convergence properties of algorithms could open up a valuable avenue for investigation.
6 Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A MISSING PROOFS AND ADDITIONAL RESULTS
A.1 A Lemma used for our First Set of Convergence Conditions
We present a lemma that is used to establish all convergence rate results for the expected loss of algorithm (1) presented in Section 3.1. The lemma relies on the following condition involving the loss function $\ell$ used by algorithm (1) and the loss function $\tilde{\ell}$ used for evaluating the performance of the algorithm.
@@ A.1 A Lemma used for our First Set of Convergence Conditions
where $\bar{\theta}_{t}=(1/t)\sum_{s=1}^{t}\theta_{s}$.
@@
A.2 Proof of Theorem 3.2
@@
A.3 Proof of Theorem 3.3
$\frac{\ell(u)}{\ell^{\prime}(u)^{2}}=\frac{1}{2}.$
$=$ $\frac{\beta}{2}\left(1-\frac{1}{1+(1/(\rho^{*}-1))(1-u)}\right)$
This condition is implied by $\pi(u)\geq\pi^{*}(\ell(u))$ where
$\pi^{*}(\ell)=\frac{\beta}{2}\left(1-\frac{1}{1+\mu\sqrt{\ell}}\right)$
with $\mu\geq\sqrt{2}/(\rho^{*}-1)$.
A.4 Proof of Theorem 3.4
@@
Proof
For $u\leq 1$, $\pi^{*}(u)=\beta(a/(a+1)-u)/(\rho-u)$, so obviously, $\pi^{*}(u)\leq\beta(1-u)/(\rho-u)$. For $0\leq u\leq 1$, we show that next that $a(1-u)/(1-u^{a})-u\leq a(1-u)$, which implies that $\pi^{*}(u)\leq(a/(a+1))\beta(1-u)/(\rho-u)$. To show the asserted inequality, by straightforward calculus it can be shown that the inequality is equivalent to $(1-(1-u))^{1-a}\geq 1-(1-a)(1-u)$ which clearly holds true.
@@
where the first inequality is by concavity of the function $1-\ell^{-1}(\ell)$ and the second inequality is by Theorem 3.4.
@@ A.6 Multi-class Classification for Linearly Separable Data
We consider margin loss functions which are according to a decreasing function of $u(x,y,\theta)$, i.e. $\ell(x,y,\theta)\equiv\ell(u(x,y,\theta))$ and ~$\tilde{\ell}(x,y,\theta)\equiv\tilde{\ell}(u(x,y,\theta))$. For example, this accomodates hinge loss function for multi-class classification Crammer and Singer (2002).
@@
Note that these conditions are equivalent to those for the binary case in (2) and (3) except for an additional factor $2$ in the first of the last above inequalities.
A.7 Implications of Conditions (2) and (3)
We discuss some implications of conditions (2) and (3). Some of this discussion will tell us about loss functions for which conditions cannot be applied.
[…]
which by integrating is equivalent to
If $\tilde{\ell}=\ell$, then
$\pi(u)=\Omega\left(\frac{1}{|u|^{2}}\right)$
~$\tilde{\ell}(u)\leq\frac{\alpha}{\beta^{2}R^{2}}\pi(u)(\rho^{*}-u)^{2}.$
which implies $\pi(u)=\Theta(1/|u|^{2})$.
A.8 Proof of Lemma 3.6
Function $\Pi$ is a convex function because, by assumption, $\pi$ is an increasing function. By (8 and Jensen’s inequality, we have
@@ A.7 Implications of Conditions (2) and (3)
A.9 Proof of Theorem 3.7
Under assumptions of the theorem, by Theorem 6.3 Bubeck (2015),
@@
$\mathbb{E}[||\theta_{t+1}-\theta^{*}||^{2}\mid x_{t},y_{t},\theta_{t}]$ $=$ $||\theta_{t}-\theta^{*}||^{2}-2\mathbb{E}[z_{t}\mid x_{t},y_{t},\theta_{t}]\nabla_{\theta}\ell_{t}(\theta_{t})^{\top}(\theta_{t}-\theta^{*})+\mathbb{E}[z_{t}^{2}\mid x_{t},y_{t},\theta_{t}]||\nabla_{\theta}\ell_{t}(\theta_{t})||^{2}$
@@
A.12 Proof of Condition (12) for $\pi(x,y,\theta)=\zeta(x,y,\theta)^{\eta}$
It suffices to show that under given conditions, (11) satisfies condition (10). Using (11), condition (10) can be written as follows
$\left(\frac{1}{2L}\right)^{\eta}\geq\frac{1}{2(1-c)}\beta^{1-\eta}.$
A.13 Sampling with Stochastic Polyak’s Step Size for Linear Binary Classifiers
@@
We next note some properties of function $h$ which allow us to establish convergence rate guarantees for some loss and uncertainty-based sampling. See Figure 5 for a graphical illustration.
@@
By Lemma A.7, the condition of Lemma 3.10 is satisfied with $\rho=1/(2L)$, by sampling proportional to absolute error loss $\pi^{*}(u)=\omega(1-\sigma(u))$ with $\beta/(4(1-c)L^{\prime})\leq\omega\leq 1$.
@@
Function $h$, defined in (13), satisifies, for every fixed $a\in(0,1/2]$,
@@
By Lemma A.8, it follows that condition of Lemma 3.10 is satisfied by uncertainty sampling according to
@@
We next consider the case $u\geq 0$. It sufficies to show that for every $u\geq 0$, $h(u)\leq h(-u)$, and hence the upper bound established for the previous case applies. The condition $h(u)\leq h(-u)$ is equivalent to
@@
A.15 Polyak’s Step Size for Multi-class Classification
@@
where function $h$ is defined in (13). Hence, condition of Lemma 3.10 holds under
@@
satisfies condition of Lemma 3.10.
@@
which is plays a key role in the condition of Lemma 3.10.
@@
A.16 Further Details on Experimental Setup
@@
A.16.1 Hyperparameter tuning
We used the Tree-structured Parzen Estimator (TPE) (Bergstra et al., 2011) algorithm in the hyperopt package (Bergstra et al., 2013) to tune the relevant hyperparameters for each method and minimise the average progressive cross entropy loss. For Polyak absloss and Polyak exponent we set the search space of $\eta$ to $[0.01,1]$ and the search space of $\rho$ to $[0,1]$. Note that the values of $\eta$ and $\rho$ influence the rate of sampling. In line with the typical goal of active learning, we aim to learn efficiently and minimise loss under some desired rate of sampling. Therefore, for every configuration of $\eta$ and $\rho$ we use binary search to find the value of $\beta$ that achieves some target empirical sampling rate. Observe that if we would not control for $\beta$, then our hyperparameter tuning setup would simply find values of $\eta$ and $\rho$ that lead to very high sampling rates, which is not in line with the goal of active learning. In the hyperparameter tuning we set the target empirical sampling rate to 50%.
A.16.2 Related to Figure 2
We similarly used binary search to find the value of $\beta$ that correspondingly achieves the two target values 1% and 10% with Polyak power function, while using the values of $\eta$ and $\rho$ that were optimised for a sampling rate of 50%. Therefore, our findings of the gains achieved for selective sampling according to stochastic Polyak’s step size are likely conservative since $\eta$ and $\rho$ were not optimised for specifically these sampling rates.
A.16.3 Related to Figures 3 and 4
We compare Polyak absolute loss sampling to absolute loss sampling and random sampling. In this setting we have no control over the sampling rate of absolute loss sampling. Hence, we first run absolute loss sampling to find an empircal sampling rate of 14.9%. We then again use binary search to find the value of $\beta$ to match this sampling rate with Polyak absolute loss sampling. Again, this setup is conservative with respect to the gains of Polyak absolute loss sampling as $\eta$ and $\rho$ were optimised for a sampling rate of 50%.
@@ A.16.3 Related to Figures 3 and 4
A.17 Additional Numerical Results
Below are some additional empirical experiments conducted on the same mushroom classification dataset (Chang and Lin, 2011) described in Section 4. The training loss as a function of the number of labeled instances that were selected for training (i.e., cost) is presented in Figure 6. Performance of different sampling methods on the hold-out testing set is presented in Figure 7 (average test loss) and Figure 8 (test accuracy). As can be seen from the figures, the relative performance of different sampling methods is consistent across different metrics.

### v3

@@ 5 Conclusion
In this paper, we have provided convergence rate guarantees for both loss and uncertainty-based active learning algorithms under various assumptions. Furthermore, we introduced the novel Adaptive-Weight Sampling (AWS) algorithm that combines sampling with an adaptive size, conforming to stochastic Polyak’s step size in expectation, and demonstrated its convergence rate guarantee, contingent on a condition related to the sampling probability function. For future research, it would be interesting to apply the methodologies presented in this paper to obtain additional convergence rate results. Additionally, exploring a theoretical study on the impact of bias and noise in the loss estimator, used for evaluating the sampling probability function, on the convergence properties of algorithms could open up a valuable avenue for investigation.
Appendix A Appendix: Proofs, Discussion, and Additional Results
A.1 Limitations & Discussion
It remains an open problem to investigate the tightness of convergence rate bounds for constant-weight sampling under the assumptions outlined in Theorem 3.3.
The convergence rate results presented in Theorems 3.3 and 3.6 pertain to smooth convex training loss functions. Future research may explore weaker assumptions regarding the training loss function.
Regarding loss-based sampling strategies, our theoretical analysis assumes an unbiased and noiseless loss estimator. Extending this to account for estimation bias and noise would be a valuable avenue for further investigation.
Our numerical results demonstrate the effectiveness of our proposed algorithm and the robustness of our theoretical findings to loss estimation bias and noise across different datasets, utilizing the logistic regression model as a binary classifier. Future work could explore the application of other classification models, such as multi-layer neural networks.
A.2 A Set of Convergence Rate Conditions
We present and prove two convergence rate lemmas for algorithm (1): the first providing sufficient conditions for a certain convergence rate and the second restricted to linear classifiers and linearly separable datasets.
The first lemma relies on the following condition involving the loss function $\ell$ used by algorithm (1) and the loss function $\tilde{\ell}$ used for evaluating the performance of the algorithm.
@@ A.2 A Set of Convergence Rate Conditions
where $\bar{\theta}_{n}=(1/n)\sum_{t=1}^{n}\theta_{t}$.
@@
For the case of linear classifiers and linearly separable datasets, it can be readily checked that Lemma A.2 implies the following lemma.
Lemma A.3.
Assume that there exist constants $\alpha,\beta>0$ such that for all $u\in\mathbb{R}$,
$\pi(u)\ell^{\prime}(u)^{2}R^{2}\leq\alpha\tilde{\ell}(u)$ (9)
and
$\pi(u)(-\ell^{\prime}(u))(\rho^{*}-u)\geq\beta\tilde{\ell}(u).$ (10)
Then, for any $\theta_{1}$ such that $||\theta_{1}-\theta^{*}||\leq S$ and $\{\theta_{t}\}_{t>1}$ according to algorithm (1) with $\gamma=\beta/\alpha$,
$\mathbb{E}\left[\sum_{t=1}^{n}\tilde{\ell}(y_{t}x_{t}^{\top}\theta_{t})\right]\leq S^{2}\frac{\alpha}{\beta^{2}}.$
Moreover, if $\tilde{\ell}$ is a convex function, then $\mathbb{E}\left[\tilde{\ell}(yx^{\top}\bar{\theta}_{n})\right]\leq S^{2}\alpha/(\beta^{2}n)$ where $\bar{\theta}_{n}=(1/n)\sum_{t=1}^{n}\theta_{t}$.
A.2.1 Discussion
We discuss some implications of conditions (9) and (10). Some of this discussion will help us identify the types of loss functions to which the conditions cannot be applied.
[…]
which by integrating is equivalent to
If $\tilde{\ell}=\ell$, then
$\pi(u)=\Omega\left(\frac{1}{|u|^{2}}\right)$
~$\tilde{\ell}(u)\leq\frac{\alpha}{\beta^{2}R^{2}}\pi(u)(\rho^{*}-u)^{2}.$
which implies $\pi(u)=\Theta(1/|u|^{2})$.
A.3 Proof of Theorem 3.1
$\frac{\ell(u)}{\ell^{\prime}(u)^{2}}=\frac{1}{2}.$
$=$ $\frac{\beta}{2}\left(1-\frac{1}{1+(1/(\rho^{*}-1))(1-u)}\right)$
This condition is implied by $\pi(u)\geq\pi^{*}(\ell(u))$ where
$\pi^{*}(\ell)=\frac{\beta}{2}\left(1-\frac{1}{1+\mu\sqrt{\ell}}\right)$
with $\mu\geq\sqrt{2}/(\rho^{*}-1)$.
A.4 Linear Classifiers: Zero-one Loss and Absolute Error Loss-based Sampling
Here we consider training loss functions satisfying:
Assumption A.4.
We consider sampling proportional to either zero-one loss or absolute error loss. The zero-one loss is defined as $\ell_{01}(u):=\mathbbm{1}_{\{u\leq 0\}}=\mathbbm{1}_{\{y\neq\mathrm{sgn}(x^{\top}\theta)\}}$. The absolute error loss is defined as $\ell_{\mathrm{abs}}(u):=2\left(\mathbbm{1}_{\{u<0\}}+\frac{1}{2}\mathbbm{1}_{\{u=0\}}\right)=|y-\mathrm{sgn}(x^{\top}\theta)|$. Sampling proportional to zero-one loss is defined as $\pi(u)=\omega\ell_{01}(u)$ for some constant $\omega\in(0,1]$, while sampling proportional to absolute error loss is defined as $\pi(u)=\omega\ell_{\mathrm{abs}}(u)$ for some constant $\omega\in(0,1/2]$.
Theorem A.5.
Assume that the loss function $\ell$ satisfies Assumption A.4 and $\rho^{*}>0$. Then, under sampling proportional to zero-one loss, for any initial value $\theta_{1}$ such that $||\theta_{1}-\theta^{*}||\leq S$ and $\{\theta_{t}\}_{t>1}$ according to algorithm (1) with $\gamma=c_{1}\rho^{*}/(c_{2}^{2}R^{2})$,
Furthermore, under sampling proportional to absolute error loss and $\gamma=2c_{1}\rho^{*}/(c_{2}^{2}R^{2})$, the bound in (11) holds but with an additional factor of $2$.
Proof is provided in Appendix A.4.1.
A.4.1 Proof of Theorem A.5
@@ A.4 Linear Classifiers: Zero-one Loss and Absolute Error Loss-based Sampling
A.5 Linear Classifiers: Generalized Smooth Hinge Loss Function
Here we consider the training loss function corresponding to the generalized smooth hinge loss function [Rennie, 2005], defined for $a\geq 1$ as follows:
$\ell(u)=\left\{\begin{array}[]{ll}\frac{a}{a+1}-u&\hbox{ if }u\leq 0\\ \frac{a}{a+1}-u+\frac{1}{a+1}u^{a+1}&\hbox{ if }0\leq u\leq 1\\ 0&\hbox{ otherwise}.\end{array}\right.$ (12)
This is a continuously differentiable function that converges to the value of the hinge loss function, $\max\{1-u,0\}$, as $a$ goes to infinity. The family of loss functions parameterized by $a$ accommodates the smooth hinge loss function with $a=1$, introduced by Rennie and Srebro [2005].
Theorem A.6.
$\pi^{*}(u)=\left\{\begin{array}[]{ll}\beta\frac{\frac{a}{a+1}-u}{\rho-u}&\hbox{ if }u\leq 0\\ \beta\frac{\frac{a}{a+1}\frac{1}{1-u^{a}}(1-u)-\frac{1}{a+1}u}{\rho-u}&\hbox{ if }0\leq u\leq 1\\ 0&\hbox{ if }u\geq 1.\end{array}\right.$
Proof is provided in Appendix A.5.1.
Note that $\pi^{*}$ is increasing in $a$ and is upper-bounded by $\pi^{**}(u)=\beta(1-u)/(\rho-1)$, which may be regarded as the limit for the hinge loss function. See Figure 3 for an illustration.
We remark that the expected loss bound in Theorem A.6 is the same as for the squared hinge loss function in Theorem 3.1, except for an additional factor $c_{a,\rho}=a/(a(\rho-1)+1)$. This factor is increasing in $a$ but always lies in the interval $[1/\rho,1/(\rho-1)]$, where the boundary values are achieved for $a=1$ and $a\rightarrow\infty$, respectively.
Theorem A.7.
Assuming $\rho^{*}>1$, the loss function is the generalized smooth hinge loss, and the sampling probability function $\pi^{*}$ satisfies the assumptions in Theorem A.6, with $n>((a+1)/a)c_{a,\rho}R^{2}S^{2}/\beta$, the expected number of sampled points is bounded as
where $\kappa$ is some positive constant.
Proof is provided in Appendix A.5.2.
For any fixed number of iterations $n$ satisfying the condition of the theorem, the expected number of sampled points is bounded by a constant for sufficiently large value of parameter $a$. The bound in Theorem A.7 depends on how the loss function $\ell(u)$ varies with $u$. When $a$ is large, $\ell(u)$ is approximately $1-u$ (hinge loss), otherwise, it is approximately $\frac{a}{2}(1-u)^{2}$ for $0\leq u\leq 1$ (squared hinge loss).
A.5.1 Proof of Theorem A.6
@@ A.5 Linear Classifiers: Generalized Smooth Hinge Loss Function
Proof.
For $u\leq 1$, $\pi^{*}(u)=\beta(a/(a+1)-u)/(\rho-u)$, so obviously, $\pi^{*}(u)\leq\beta(1-u)/(\rho-u)$. For $0\leq u\leq 1$, we show that next that $a(1-u)/(1-u^{a})-u\leq a(1-u)$, which implies that $\pi^{*}(u)\leq(a/(a+1))\beta(1-u)/(\rho-u)$. To show the asserted inequality, by straightforward calculus it can be shown that the inequality is equivalent to $(1-(1-u))^{1-a}\geq 1-(1-a)(1-u)$ which clearly holds true. ∎
@@
where the first inequality is by concavity of the function $1-\ell^{-1}(\ell)$ and the second inequality is by Theorem A.6.
@@ A.6 Multi-class Classification for Linearly Separable Data
We consider margin loss functions which are according to a decreasing function of $u(x,y,\theta)$, i.e. $\ell(x,y,\theta)\equiv\ell(u(x,y,\theta))$ and ~$\tilde{\ell}(x,y,\theta)\equiv\tilde{\ell}(u(x,y,\theta))$. For example, this accomodates hinge loss function for multi-class classification Crammer and Singer [2002].
@@
Note that these conditions are equivalent to those for the binary case in (9) and (10) except for an additional factor $2$ in the first of the last above inequalities. ∎
A.7 Proof of Lemma 3.2
Function $\Pi$ is a convex function because, by assumption, $\pi$ is an increasing function. By (4 and Jensen’s inequality, we have
@@
A.8 Proof of Theorem 3.3
Under assumptions of the theorem, by Theorem 6.3 Bubeck [2015],
@@
$\mathbb{E}[||\theta_{t+1}-\theta^{*}||^{2}\mid x_{t},y_{t},\theta_{t}]$
$=$ $||\theta_{t}-\theta^{*}||^{2}-2\mathbb{E}[z_{t}\mid x_{t},y_{t},\theta_{t}]\nabla_{\theta}\ell_{t}(\theta_{t})^{\top}(\theta_{t}-\theta^{*})+\mathbb{E}[z_{t}^{2}\mid x_{t},y_{t},\theta_{t}]||\nabla_{\theta}\ell_{t}(\theta_{t})||^{2}$
@@
A.11 Proofs of Corollaries 3.7 and 3.8
@@
A.11.1 Proof of Corollary 3.7
We first note the following lemma, whose proof is provided in Appendix A.12.
@@
See See Figure 4, left, for a graphical illustration.
By Lemma A.12, condition (6) in Theorem 3.6 is satisfied with $\rho=1/(2L)$, by sampling proportional to absolute error loss $\pi^{*}(u)=\omega(1-\sigma(u))$ with $\beta/(4(1-c)L^{\prime})\leq\omega\leq 1$.
@@
A.11.2 Proof of Corollary 3.8
We have the following lemma, whose proof is provided in A.13.
@@
Function $h$, defined in (20), satisfies, for every fixed $a\in(0,1/2]$,
@@
See See Figure 4, right, for a graphical illustration.
By Lemma A.13, it follows that condition (6) in Theorem 3.6 is satisfied by uncertainty sampling according to
@@
We next consider the case $u\geq 0$. It suffices to show that for every $u\geq 0$, $h(u)\leq h(-u)$, and hence the upper bound established for the previous case applies. The condition $h(u)\leq h(-u)$ is equivalent to
@@
A.14 Convergence Conditions for $\pi(x,y,\theta)=\zeta(x,y,\theta)^{\eta}$
It suffices to show that under given conditions, the sampling probability function satisfies condition (6). Using the definition of the sampling probability function, condition (6) can be written as follows
$\left(\frac{1}{2L}\right)^{\eta}\geq\frac{1}{2(1-c)}\beta^{1-\eta}.$
A.15 Uncertainty-based Sampling for Multi-class Classification
@@
where function $h$ is defined in (20). Hence, condition of Theorem 3.6 holds under
@@
satisfies condition of Theorem 3.6.
@@
which is plays a key role in the condition of Theorem 3.6.
@@
Appendix B Appendix: Additional Material for Numerical Experiments
B.1 Further Details on Experimental Setup
Hyperparameter Tuning
We used the Tree-structured Parzen Estimator (TPE) [Bergstra et al., 2011] algorithm in the hyperopt package [Bergstra et al., 2013] to tune the relevant hyperparameters for each method and minimize the average progressive cross entropy loss. For Polyak absloss and Polyak exponent we set the search space of $\eta$ to $[0.01,1]$ and the search space of $\rho$ to $[0,1]$. Note that the values of $\eta$ and $\rho$ influence the rate of sampling.
In line with the typical goal of active learning, we aim to learn efficiently and minimize loss under some desired rate of sampling. Therefore, for every configuration of $\eta$ and $\rho$ we use binary search to find the value of $\beta$ that achieves some target empirical sampling rate.
Observe that if we would not control for $\beta$, then our hyperparameter tuning setup would simply find values of $\eta$ and $\rho$ that lead to very high sampling rates, which is not in line with the goal of active learning. In the hyperparameter tuning we set the target empirical sampling rate to 50%.
Compute Resources
All experiments were performed on a single machine with 72 CPU cores and 228 GB RAM. It took us around 2,000 seconds to complete a training run for an AWS-PA with an absloss estimator on Mushrooms dataset, our slowest experiment. The training runs for other datasets and algorithms were considerably faster.
B.2 Further Details on Numerical Experiments with Different Algorithms
In Section 4, we presented numerical results for comparing AWS-PA with other algorithms. These results are shown in Figure 1. Below are some additional details for these experiments.
Tuning Sampling Rate
In Figure 1 we compare Polyak absolute loss sampling to absolute loss sampling and random sampling. In this setting we have no control over the sampling rate of absolute loss sampling. Hence, we first run absolute loss sampling to find an empirical sampling rate of 14.9%. We then again use binary search to find the value of $\beta$ to match this sampling rate with Polyak absolute loss sampling. Again, this setup is conservative with respect to the gains of Polyak absolute loss sampling as $\eta$ and $\rho$ were optimized for a sampling rate of 50%.
Sampling Efficiency of AWS-PA
In Figure 1 we had demonstrated on various datasets that AWS-PA leads to faster convergence than the traditional loss-based sampling Yoo and Kweon [2019]. Figure 5 presents results as a function of the number of sampled instances, i.e., the number of labeled instances that were selected for training (i.e., cost). This contrasts Figure 1, which showed on the X-axis the total number of iterations. The results confirm that sampling with AWS-PA not only leads to faster convergence than traditional loss-based sampling when expressed in terms of number of iterations, but also when expressed in the number of sampled instances.
@@ B.2 Further Details on Numerical Experiments with Different Algorithms
AWS-PA Results on a Holdout Test Set
The results in Figure 1 were obtained using a progressive validation Blum et al. [1999] procedure where the average loss is measured during an online learning procedure where for each instance the loss is calculated prior to the weight update. Figures 6 and 7 show that our finding that AWS-PA leads to faster convergence than traditional loss-based sampling and than random sampling also holds true on a separate hold out test set.
@@
B.3 Further Details on the Robustness of AWS-PA to Loss Estimation
In Section 4, we presented numerical results for comparison of the training loss achieved by our AWS-PA algorithm using the ground truth absolute error loss and estimated absolute error loss. These results are shown in Figure 2. Here, we provide more details for the underlying setup of experiments, and the number of sampled points.
Details on the Absloss Estimator
For the experiments in Figure 2 we use a separate Random Forest (RF) regressor which estimates absolute error loss based on the same set of features as the target model with an addition of the target’s model prediction as an extra feature. The estimator is retrained on every sampling step using the labeled points observed so far. We used the scikit-learn implementation of the RF regressor and manually tuned two hyperparameters for different datasets: (a) number of tree estimators (b) number of "warm-up" steps during which we sample content with a constant probability until we collect enough samples to train an RF estimator. We parallelized training of the RF estimator across all available CPU cores and used default values for all other hyperparameters.
The statistics of the absloss as well as the parameters of the RF estimators for different datasets are summarized in Table 2. From the table, we note that the mean ground truth values of the absloss are largely in line with the mean estimated absloss. This suggests that it is possible to train absloss estimator with low bias or is even unbiased.
Sampling efficiency of the absloss estimator
In Figure 8 we compare the cost of sampling based on the ground truth absolute loss versus sampling based on the estimated absloss. We note that in 4 out 6 datasets, the sampling cost closely matches that of sampling based on the ground truth absloss. However, in one of the cases (Splice) the sampling cost is lower and in one of the cases (Credit) it is higher than the baseline.
Table 2: Hyperparameters of the absloss estimator and the comparison of the mean of ground truth and the mean of estimated absolute loss values.
Dataset Number of trees Warm-up steps Mean absloss Mean estimated absloss
Mushrooms 25 1 0.100 0.104
MNIST 3s vs 5s 25 50 0.089 0.087
Parkinsons 100 5 0.448 0.443
Splice 100 25 0.163 0.160
Tictactoe 100 1 0.435 0.416
Credit 100 50 0.277 0.294
(a) Mushrooms dataset
(b) MNIST classification of 3s vs 5s
(c) Parkinsons dataset
(d) Splice dataset
(e) Tictactoe dataset
(f) Credit dataset
Figure 8: Sampling efficiency for sampling based on the ground truth of absolute loss v.s. on estimated absolute loss.
B.4 Additional experiments
Experiments with different sampling rates
Loizou et al. [2021] demonstrated that stochastic gradient descent with a step size corresponding to their stochastic Polyak’s step size converges faster than gradient descent. Figure 9 illustrates that these findings extend to scenarios where we selectively sample from the dataset rather than training on the full dataset, and the step size is according to stochastic Polyak’s step size only in expectation.
Figure 9: Average cross-entropy progressive loss of Polyak’s step size compared to SGD with constant step size, for 1% and 10% sampling from the mushrooms data.
To perform these experiments, similarly to the procedure described in Appendix B.1, we used binary search to find the value of $\beta$ that correspondingly achieves the two target values 1% and 10% with Polyak power function, while using the values of $\eta$ and $\rho$ that were optimised for a sampling rate of 50%. Therefore, our findings of the gains achieved for selective sampling according to stochastic Polyak’s step size are likely conservative since $\eta$ and $\rho$ were not optimised for specifically these sampling rates.
Experiments with synthetic absloss estimator
We simulate a noisy estimator of the absolute error loss in AWS-PA. We model an unbiased noisy estimator ^$\hat{\ell}_{\mathrm{abs}}$ of the absolute error loss $\ell_{\mathrm{abs}}\in[0,1]$ as a random variable following the beta distribution, denoted as ^$\hat{\ell}_{\mathrm{abs}}\sim\mathrm{Beta}(\alpha,\beta)$, where $\alpha$ and $\beta$ are parameters set to ensure $\mathbb{E}[\hat{\ell}_{\mathrm{abs}}]=\ell_{\mathrm{abs}}$. The variance of the noise can be controlled by the tuning parameter $\alpha$, given by
$\mathrm{var}[\hat{\ell}_{\mathrm{abs}}]=\frac{\ell_{\mathrm{abs}}(1-\ell_{\mathrm{abs}})}{\alpha+\ell_{\mathrm{abs}}}.$
Figure 10 shows that the convergence results are robust against estimation noise of the absolute error loss for a wide range of values for $\alpha\geq 1$.